[Hallc_running] [New Logentry] RC daily update

hszumila at jlab.org hszumila at jlab.org
Tue Aug 11 09:25:02 EDT 2020

Logentry Text:
Hall is currently in Controlled Access for the scheduled work during the RF recovery. 

We finished our BCM calibration yesterday with max current of 34 uA. Dave Mack is looking into the runs. 
Junhao is setting target from 270->90 deg and will do an EPR measurement. 
Joe B. will do a walk through this morning. 

We had some alarms going off throughout owl for a noisy readback on the charge asymmetry after changing the IHWP, but Hall A says there is no issue. Looking into this now with Ops. https://logbooks.jlab.org/entry/3824780 

I think in total, we currently have around 3.5-4 shifts of beam on 3He production target. 
Murchana and Bill will do a Moller measurement tomorrow during the Day shift with Dave G. monitoring. 

We meet today at 4:15pm via bluejeans (861439920): https://hallcweb.jlab.org/wiki/index.php/RC_Update_and_Run_Plan_for_Aug_11,_2020


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3824842
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200811/d5b1b1d7/attachment.html>

More information about the Hallc_running mailing list